what is the slope and y-intercept of 4x+5y=-20
Answer by tw23279(14)

Move the equation around! That's the heart of Algebra!
So since 4x + 5y = -20
5y = -4x - 20
y= (-4x - 20) / 5
Since -4x and 20 are TERMS, not factors. we have to divide them both individually by 5, and we get that:
y= (-4/5x) - 4
And we know that this is in the form y = mx + b
Just look at it
And m = slope = -4/5
b = y-intercept = -4